# Tuning-Free Alignment of Diffusion Models with Direct Noise Optimization

### v1

@@ 1 Introduction
In recent years, diffusion models have become the state-of-the-art model for generating high-quality images, with exceptional resolution, fidelity, and diversity (Ho et al., 2020; Dhariwal and Nichol, 2021; Song et al., 2020b). These models are also remarkably easy to train and can be effectively extended to conditional generation (Ho and Salimans, 2022). Specifically, diffusion models work by learning to reverse the process of diffusing the data distribution $p(x)$ into noise, which can be described by a stochastic differential equation (SDE) (Song et al., 2020b; Karras et al., 2022): $d{x}_{t}=f(t){x}_{t}dt+g(t)dw_{t},$ where $dw_{t}$ is the standard Wiener process, and $f(t)$ and $g(t)$ are the drift and diffusion coefficients, respectively. The reverse process relies on the score function $\epsilon({x}_{t},t)\stackrel{{\scriptstyle\text{def.}}}{{=}}\nabla_{x}\log p_{t}({x})$ where $p_{t}$ denotes the p.d.f of noisy data $x_{t}$, and its closed-form can be expressed either as an ODE or as an SDE: (Song et al., 2020b):
@@
Despite the effectiveness of diffusion models in modeling continuous distributions, a prevalent issue of diffusion model is that they are trained on large-scale, unlabeled, and mixed-quality data to capture the data distribution. However, when deploying these generative models for specific tasks, it is no longer suitable to sample from the original learned distribution directly, because this distribution has not been aligned with the task-specific objective. For instance, in image generation, users aim to generate images that are aesthetically pleasing rather than random ones. Hence, it is imperative to tailor the generative models to align with the task-specific objectives before utilization. The topic of aligning generative models has gained significant attention following the success of aligning LLMs with Reinforcement Learning from Human Feedback (RLHF) (Ouyang et al., 2022; Liu et al., 2023a; OpenAI, 2022; Bai et al., 2022). Recently, the alignment problem has also drawn considerable interest in the context of diffusion models, as evidenced by a series of studies such as (Deng et al., 2024; Uehara et al., 2024; Yuan et al., 2024; Song et al., 2023a; Dong et al., 2023; Sun et al., 2023; Zhang et al., 2023; Prabhudesai et al., 2023; Black et al., 2023; Fan et al., 2023).
Alignment Problem for Diffusion Models. Given a diffusion model characterized by parameters $\theta$ and its associated distribution $p_{\theta}(x)$, as well as a reward function $r(x)$ that can assign real-valued scores to generated samples, the central goal of the alignment problem is to adjust the distribution $p_{\theta}(x)$ such that the generated samples maximize the reward from $r(x)$. Such reward models $r(x)$ can be constructed by training a neural network on human feedback data (Kirstain et al., 2023) or by utilizing custom-designed metrics. In this work, we consider the reward models to be continuous and possibly non-differentiable. In the following sections, we will provide a comprehensive review of some well-established methods for aligning diffusion models.
@@ 1.1.1 Online/Offline Reinforcement Learning for Fine-tuning Diffusion Models
Aligning generative models with specific objectives, such as human preferences, has attracted extensive research efforts, following a series of successful attempts that aligns LLMs with Reinforcement Learning from Human Feedback (RLHF). For diffusion models, developing alignment algorithms has also gained sufficient attention recently, where a majority of recent research has focused on leveraging Reinforcement Learning (RL)-based approaches. A common mathematical formulation in RL-based methods is to maximize the expected reward while ensuring the resulting distribution does not deviate excessively from the original distribution. This can be expressed as the following KL-regularized optimization problem:
@@ 1.1.2 Direct Fine-tuning of Diffusion Models with Differentiable Rewards
Solving the ODE (1) or the SDE (2) typically involves discretizing the time steps into $T$ steps. By starting with the initial noise $x_{T}\sim\mathcal{N}(0,I)$, the solution process for the ODE (1) can be viewed as a mapping that transforms the initial noise $x_{T}$ into less noisy data through
$x_{t-1}=\text{ODE\_solver}(x_{t}),$ (4)
for $t=T,...,1$, eventually resulting in the generated sample $x_{0}$. Similarly, solving the SDE (2) can be seen as a mapping that gradually converts both the initial noise $x_{T}$ and the entire discretized Wiener process $\{w_{t}\}$ into the generated sample $x_{0}$, e.g., through
$x_{t-1}=\text{SDE\_solver}(x_{t})+w_{t}.$ (5)
Diffusion Sampling as a Noise-to-Sample Mapping. From this viewpoint, the diffusion sampling process can be conceptualized as an end-to-end mapping $M_{\theta}(z)$, which translates noise vectors $z$, sampled from the standard Gaussian distribution, into generated samples. Here, the noise vectors $z$ serve as a unified abstraction for both the $x_{T}$ in the ODE-based sampling process and the $(x_{T},\{w_{t}\})$ in the SDE-based sampling process.
Two recent studies, AlignProp (Prabhudesai et al., 2023) and DRaFT (Clark et al., 2023), consider directly fine-tune diffusion models using differentiable rewards. Specifically, their optimization objective is formulated as:
@@ 1.2 Comparing Existing Methods
Existing works can generally be categorized based on two criteria: whether it requires fine-tuning and whether the reward model needs to be differentiable.
Tuning-based or Tuning-free. All RL-based methods and the direct fine-tuning method are tuning-based, meaning they necessitate adjustments to the network models $\theta$. There are two main disadvantages associated with tuning-based methods. The first is that they require fine-tuning for each new reward model, a process that can consume considerable resources, especially when faced with extensive choices for reward models. The second drawback is that the fine-tuning process typically relies on a limited set of input prompts, which challenges the model to generalize to new and unseen prompts. In contrast, methods such as LGD (Song et al., 2023a) belong to the tuning-free category. The main advantage of this approach is its elimination of the need for fine-tuning, allowing it to be flexibly applied in conjunction with any reward models and input prompts. Further, tuning-free methods require significantly fewer computing resources than the tuning-based methods. However, the major drawback of tuning-free methods is the substantial increase in the time required for the generation process. Also, our experiments in Section 5.2 demonstrate that the tuning-free method LGD cannot match the performance of tuning-based method.
Differentiable or Non-Differentiable. Current RL-based methods can work by utilizing solely the value or preference information of the reward functions. In contrast, the existing direct fine-tuning methods and LGD require the reward function to be differentiable.
@@ 1.3 Our Contributions
In this work, we concentrate on tuning-free alignment of diffusion models, as we believe that flexibility and low computing requirements matter more than generation time for a broad range of real-world applications. Our primary goal is to design a tuning-free method that can achieve the same level or even better performance as tuning-based methods, and is capable of handling both differentiable and non-differentiable objective functions. To this end, we focus on an under-explored route for achieving tuning-free alignment of diffusion models—Direct Noise Optimization (DNO). Specifically, we make the following contributions:
@@
  • We identify out-of-distribution reward-hacking as a critical issue in DNO. To address this problem, we introduce a probability-regularized noise optimization method designed to ensure the generated samples remain within the support of pretrained distribution.
  • By devising a novel and highly efficient hybrid gradient approximation strategy, we extend the DNO approach to handle non-differentiable reward functions effectively.
@@ 2 Direct Noise Optimization for Aligning Diffusion Models
$\max_{z}r(M_{\theta}(z)).$ (7)
By solving this optimization problem, we can obtain the optimized noise vectors, which are then used to generate refined samples. When the reward function $r(\cdot)$ is differentiable, gradient-based optimization methods can be applied to solve the problem efficiently. As one can see, this method is tuning-free, similar to LGD Song et al. (2023a).
@@
Although idea of DNO have appeared in existing works, many aspects of it still remain under-explored. On one hand, there is a lack of comprehensive framework concerning the design choices, theoretical understanding, and practical challenges associated with DNO for aligning diffusion models. On the other hand, the field has yet to systematically investigate whether DNO, as a tuning-free method for aligning diffusion models, can achieve competitive performance of tuning-based methods. In this work, we aim to conduct a thorough study on DNO.
In the following two sections, we dive deep to understand the theoretical foundation of DNO and discuss a critical design choice.
@@
Figure 2: Evolution of optimized distribution.
2.1 Understanding Noise Optimization
First, we present a simple example to visualize the process of noise optimization. Specifically, we trained a toy diffusion model for the uniform distribution on a ring with a radius between $0.8$ and $1.2$, and the initial distribution is visualized in Figure 2(a), where each red point denotes a single sample drawn from the trained diffusion model. Then, we solve the noise optimization problem (7) with the reward function $r(x)=\sin(4\pi x[1])+\sin(4\pi x[2])-\left((x[1]-1)^{2}-x[2]^{2}\right)/5$, a highly nonconvex function with many local maxima. In Figures 2(b), 2(c), and 2(d), we visualize the optimized samples after 10, 50, and 100 gradient steps, respectively. We can observe that the distribution of the samples shifts toward a distribution on the local maxima of the reward function.
Noise Optimization as Sampling from an Improved Distribution. As evident from Figure 2, optimizing the noise vectors for diffusion models indeed leads to an improved distribution in terms of the reward scores. To rigorously describe this evolving process, we define an operator function $g$ to represent a single gradient step, i.e., $g(z)\stackrel{{\scriptstyle\text{def}}}{{=}}z+\ell\nabla_{z}r\circ M_{\theta}(z)$, where $\ell$ denotes the step size for gradient ascent. Additionally, we define the operator $g_{t}$, which denotes applying the gradient ascent step for $t$ steps, i.e., $g_{t}(z)=g(g_{t-1}(z))$ with $g_{0}$ being the identity mapping. With these notations, we can now express the distribution after $t$ gradient steps as $p_{t}(x)$, which is the distribution of $M_{\theta}(g_{t}(z))$ with $z\sim\mathcal{N}(0,I)$. In the following theorem, we demonstrate that there is a rigorous improvement after every single gradient step, i.e., the distribution $p_{t+1}(x)$ is provably better than $p_{t}(x)$ in terms of expected reward.
@@ 2.1 Understanding Noise Optimization
As depicted in (8), provided that the step size $\ell$ is less than $\frac{2}{L}$, the distribution $p_{t+1}(x)$ is strictly better than the previous distribution $p_{t}(x)$, as long as the second term is non-zero. Besides, a natural question arises: When does the distribution stop improving? Namely, when does the second term in (8) become zero. To answer this question, we provide a detailed discussion on this point and the proof for Theorem 1 in Appendix B.
@@
As previously introduced, there are two primary methods for sampling from pretrained diffusion models: one based on solving the ODE (1) and the other on solving the SDE (2). A critical difference lies in that ODE sampling depends exclusively on the initial noise $x_{T}$, whereas SDE sampling is additionally influenced by the noise ${w_{t}}$ added at every step of the generation process. It has been noted that existing works on noise optimization (Ben-Hamu et al., 2024; Novack et al., 2024; Karunratanakul et al., 2023) have mainly concentrated on optimizing the initial noise $x_{T}$ for the ODE sampler.
Figure 3: ODE vs. SDE for optimization
Different from preceding studies, we explore the utilization of the SDE sampler for noise optimization. Specifically, we employ the DDIM with $\eta=1$ (Song et al., 2020a) as the SDE sampler and propose to optimize both the added noise $w_{t}$ at every timestep and the initial noise $x_{T}$. In this context, the dimensionality of the optimized ßnoise significantly surpasses that of ODE sampler. Empirically, we have observed that optimizing the SDE-based generation process can be significantly faster than its ODE-based counterpart. This is illustrated in Figure 3, where we juxtapose the optimization speeds between the ODE (DDIM with $\eta=0$) and SDE samplers using both a toy example described above (Left) and optimizing for stable diffusion in alignment with the aesthetic reward (Right), a major experiment detailed in the subsequent section 5.2. Intuitively, this acceleration in optimization speed can be attributed to the finer-grained control over the generation process afforded by the SDE sampler. Given these observations, our work will concentrate on optimizing the SDE sampler for the remainder of the study, and fix the generation steps to be 50.

### v2

@@ 1 Introduction
Figure 2: DNO performs alignment without any modification to the diffusion models. $M_{\theta}$ denotes the sampling process of diffusion models, which maps the noise distribution into the sample distribution. Existing methods for aligning diffusion models seek to modify the noise-to-sample mapping $M_{\theta}$ through either fine-tuning network $\theta$ or gradient-based guidance. In contrast, DNO does not change $M_{\theta}$; instead, it directly optimizes the noise vectors for alignment during inference.
In recent years, diffusion models have become the state-of-the-art model for generating high-quality images, with exceptional resolution, fidelity, and diversity (Ho et al., 2020; Dhariwal and Nichol, 2021; Song et al., 2020b). These models are also remarkably easy to train and can be effectively extended to conditional generation (Ho and Salimans, 2022). Specifically, diffusion models work by learning to reverse the process of diffusing the data distribution $p(x)$ into noise, which can be described by a stochastic differential equation (SDE) (Song et al., 2020b; Karras et al., 2022):
$d{x}_{t}=f(t){x}_{t}dt+g(t)dw_{t},$ (1)
where $dw_{t}$ is the standard Wiener process, and $f(t)$ and $g(t)$ are the drift and diffusion coefficients, respectively. The reverse process relies on the score function $\epsilon({x}_{t},t)\stackrel{{\scriptstyle\text{def.}}}{{=}}\nabla_{x}\log p_{t}({x})$ where $p_{t}$ denotes the p.d.f of noisy data $x_{t}$, and its closed-form can be expressed either as an ODE or as an SDE: (Song et al., 2020b):
@@
Despite the effectiveness of diffusion models in modeling continuous distributions, a prevalent issue of diffusion models is that they are trained on large-scale, unlabeled, and mixed-quality data to capture the data distribution. However, when deploying these generative models for specific tasks, it is no longer suitable to sample from the original learned distribution directly, because this distribution has not been aligned with the task-specific objective. For instance, in image generation, users aim to generate images that are aesthetically pleasing rather than random ones. Hence, it is imperative to tailor the generative models to align with the task-specific objectives before deployment. The topic of aligning generative models has gained significant attention following the success of aligning LLMs with Reinforcement Learning from Human Feedback (RLHF) (Ouyang et al., 2022; Liu et al., 2023a; OpenAI, 2022; Bai et al., 2022). Recently, the alignment problem has also drawn considerable interest in the context of diffusion models, as evidenced by a series of studies such as (Deng et al., 2024; Uehara et al., 2024; Yuan et al., 2024; Song et al., 2023a; Dong et al., 2023; Sun et al., 2023; Zhang et al., 2023; Prabhudesai et al., 2023; Black et al., 2023; Fan et al., 2023).
Alignment Problem for Diffusion Models. Given a diffusion model characterized by parameters $\theta$ and its associated distribution $p_{\theta}(x)$, as well as a reward function $r(x)$ that can assign real-valued scores to generated samples, the central goal of the alignment problem is to adjust the distribution $p_{\theta}(x)$ such that the generated samples maximize the reward from $r(x)$. Such reward models $r(x)$ can be constructed by training a neural network on human feedback data (Kirstain et al., 2023) or by utilizing custom-designed metrics. In this work, we consider the reward models to be continuous but possibly non-differentiable. In the following sections, we will provide a comprehensive review of some well-established methods for aligning diffusion models.
@@ 1.1.1 Online/Offline Reinforcement Learning for Fine-tuning Diffusion Models
Aligning generative models with specific objectives, such as human preferences, has attracted extensive research efforts, following a series of successful attempts that align LLMs with Reinforcement Learning from Human Feedback (RLHF). For diffusion models, developing alignment algorithms has also gained sufficient attention recently, where a majority of recent research has focused on leveraging Reinforcement Learning (RL)-based approaches. A common mathematical formulation in RL-based methods is to maximize the expected reward while ensuring the resulting distribution does not deviate excessively from the original distribution. This can be expressed as the following KL-regularized optimization problem:
@@ 1.1.2 Direct Fine-tuning of Diffusion Models with Differentiable Rewards
Before delving into the formal description of this method, it is useful to revisit the sampling process of diffusion models, which also serves as the foundation for the rest of this work. Solving the ODE (2) or the SDE (3) typically involves discretizing the time steps into $T$ steps. By starting with the initial noise $x_{T}\sim\mathcal{N}(0,I)$, the solution process for the ODE (2) can be viewed as a mapping that transforms the initial noise $x_{T}$ into less noisy data through the following process:
$x_{t-1}=\text{ODE\_solver}(x_{t}),\text{ for }t=T,...,1.$ (5)
After $T$ steps, the output will be the generated sample $x_{0}$. Similarly, solving the SDE (3) can be seen as a mapping that gradually converts both the initial noise $x_{T}$ and the entire extra random noise vectors $z_{T},...,z_{1}$ into the generated sample $x_{0}$, e.g., through the following process:
$x_{t-1}=\text{SDE\_solver}(x_{t},z_{t}),\text{ for }t=T,...,1,$ (6)
where $z_{t}$ is also drawn from standard Gaussian distribution. In the example below, we provide a specific example of how the above process works in practice.
Example. In Algorithm 1 described below, we provide an example of the sampling algorithm for diffusion models, DDIM (Song et al., 2020a), which is essentially the Euler method for solving ODEs/SDEs. The diffusion coefficients $\alpha_{1},\ldots,\alpha_{T}$ in Algorithm 1 are computed using the coefficient functions $f(t)$ and $g(t)$ in (1), with the detailed computation found in (Song et al., 2020b; Karras et al., 2022). The coefficient $\eta$ in DDIM determines whether we are solving the ODE or the SDE, with $\eta=0$ corresponding to ODE and $\eta>0$ corresponding to SDE.
0:  Discretization timesteps $T$, diffusion coefficient $\alpha_{1},...,\alpha_{t}$, initial noise $x_{T}\sim{\mathcal{N}}(0,I)$, noise vectors $z_{1},...,z_{T}\sim{\mathcal{N}}(0,I)$, coefficient $\eta\in[0,1]$ for balancing ODE and SDE, learned score network $\epsilon_{\theta}(\cdot,\cdot)$.
1:  for $t=T$ to $1$ do
2:     Compute $\sigma_{t}=\eta\sqrt{(1-\alpha_{t-1}/(1-\alpha_{t})}\sqrt{1-\alpha_{t}/\alpha_{t-1}}$.
3:     $x_{t-1}=\sqrt{\alpha_{t-1}/\alpha_{t}}\cdot x_{t}-\left(\sqrt{\alpha_{t-1}(1-\alpha_{t})/\alpha_{t}}-\sqrt{1-\alpha_{t-1}-\sigma_{t}^{2}}\right)\epsilon_{\theta}(x_{t},t)+\sigma_{t}z_{t}$
4:  end for
5:  return  $x_{0}$
Algorithm 1 DDIM Sampling Algorithm
Remark 1.
Throughout this work, for simplicity, we only adopt the DDIM sampling algorithm for our experiments, as it remains one of the most popular choices for diffusion sampling and, more importantly, supports both ODE-style and SDE-style sampling.
Diffusion Sampling as a Noise-to-Sample Mapping. From the diffusion sampling process described in (5) and (6), we observe that the sampling process can be conceptualized as an end-to-end mapping $M_{\theta}(z)$, which translates noise vectors $z$, sampled from the standard Gaussian distribution, into generated samples. Here, the noise vectors $z$ serve as a unified abstraction for both the $x_{T}$ in the ODE-based sampling process and the $(x_{T},\{z_{1},...,z_{T}\})$ in the SDE-based sampling process. As we can see, the noise vector $z$ uniquely determines the generated sample from the diffusion models.
Two recent studies, AlignProp (Prabhudesai et al., 2023) and DRaFT (Clark et al., 2023), have proposed to directly fine-tune diffusion models using differentiable rewards. Specifically, their optimization objective is formulated as:
@@
Both the AlignProp (Prabhudesai et al., 2023) and DRaFT (Clark et al., 2023) methods utilize the ODE-type DDIM solver for the sampling process, specifically employing Algorithm 1 with $\eta=0$.
@@ 1.2 Comparing Existing Methods
Existing works can be generally categorized based on two criteria: whether it requires fine-tuning and whether the reward model needs to be differentiable.
Tuning-based or Tuning-free Methods. All RL-based methods and the direct fine-tuning method are tuning-based, meaning they necessitate adjustments to the network models $\theta$. There are two main disadvantages associated with tuning-based methods. The first is that they require fine-tuning for each new reward model, a process that can consume considerable resources, especially when faced with extensive choices for reward models. The second drawback is that the fine-tuning process typically relies on a limited set of input prompts, which challenges the model to generalize to new and unseen prompts. In contrast, methods such as LGD (Song et al., 2023a) belong to the tuning-free category. The main advantage of this approach is its elimination of the need for fine-tuning, allowing it to be flexibly applied in conjunction with any reward models and input prompts. Further, tuning-free methods require significantly fewer computing resources than tuning-based methods. However, the major drawback of tuning-free methods is the substantial increase in the time required for the generation process. Also, our experiments in Section 5.2 demonstrate that the tuning-free method LGD cannot match the performance of tuning-based methods such as AlignProp (Prabhudesai et al., 2023) and DDPO (Karunratanakul et al., 2023).
Differentiable or Non-Differentiable Rewards. Current RL-based methods can work by utilizing solely the value or preference information of the reward functions, therefore they can still work if the reward function is non-differentiable. In contrast, the existing direct fine-tuning methods and LGD require the reward function to be differentiable. In practice, working with non-differentiable reward functions is important due to their prevalence. This non-differentiable property can arise from the simulation-based procedures used to compute the reward, or the reward function itself may be a black box, so one cannot directly obtain the gradient information.
@@ 1.3 Our Contributions
In this work, we focus on tuning-free alignment of diffusion models, as we believe that flexibility and low computing requirements are critical for a broad range of real-world applications. Our primary goal is to design a tuning-free method that can achieve or outperform tuning-based methods, and is capable of handling both differentiable and non-differentiable objective functions. To this end, we focus on an under-explored technique for achieving tuning-free alignment of diffusion models—Direct Noise Optimization (DNO). Specifically, we make the following contributions:
@@
  • We identify out-of-distribution reward-hacking as a critical issue in DNO. To address this issue, we introduce a novel probability-regularized noise optimization method designed to ensure the generated samples remain within the support of pretrained distribution.
  • By developing a novel and highly efficient hybrid gradient approximation strategy, we extend the DNO approach to handle non-differentiable reward functions effectively.
@@ 2 Direct Noise Optimization for Aligning Diffusion Models
$\max_{z}\;r(M_{\theta}(z)),$ (8)
with $z\sim{\mathcal{N}}(0,I)$ as the initial solution. As we will discuss in Section 3, the Gaussian distribution ${\mathcal{N}}(0,I)$ serves as an important prior on the optimization variables $z$. By solving this optimization problem, we can obtain the optimized noise vectors, which are then used to generate refined samples.
We highlight a prominent difference between the DNO in (8) and major existing alignment methods, including LGD (Song et al., 2023a), DDPO (Black et al., 2023), and AlignProp (Prabhudesai et al., 2023). The core idea of these existing alignment methods is to modify the noise-to-sample mapping $M_{\theta}$ into a new mapping (say $M_{\text{new}}$) such that it can map the Gaussian noise distribution into high-reward samples. The tuning-free method LGD modifies $M_{\theta}$ by augmenting the diffusion ODE/SDE with reward gradient, while the tuning-based methods DDPO and AlignProp modify $M_{\theta}$ by tuning the network parameter $\theta$. Due to this observation, we refer to these existing alignment methods as Alignment Methods with Mapping-Optimization. In sharp contrast, DNO does not seek to modify the mapping $M_{\theta}$. Instead, it aims to find the sub-area of the noise distribution that can be mapped to high-reward samples. See Figure 2 for a visual illustration of this point.
When the reward function $r(\cdot)$ is differentiable, gradient-based optimization methods can be applied to solve the problem efficiently. That is, the following step can be performed iteratively until convergence:
$z_{\text{new}}=\text{optimizer\_step}(z_{\text{old}},\nabla_{z}r\left(M_{\theta}(z_{\text{old}})\right)),$ (9)
where the optimizer can be either vanilla gradient ascent or adaptive optimization algorithms like Adam (Kingma and Ba, 2014). When the gradient of reward function $r(\cdot)$ is not available, we can leverage techniques from zeroth order optimization (Nesterov and Spokoiny, 2017; Tang et al., 2024a) to estimate the ground-truth gradient $\nabla_{z}r\left(M_{\theta}(z_{\text{old}})\right)$, denoted as $\hat{g}(z_{\text{old}})$, and then apply similarly
$z_{\text{new}}=\text{optimizer\_step}(z_{\text{old}},\hat{g}(z_{\text{old}})).$ (10)
In Section 4, we will reveal that, to efficiently handle non-differentiable reward functions, the best practice for obtaining $\hat{g}(z_{\text{old}})$ is to estimate only the gradient $\nabla_{x}r(x)$, while computing the ground-truth gradient for $\nabla_{z}M_{\theta}(z)$.
We provide a visual illustration in Figure 3 to describe the main procedure for DNO using the DDIM sampling procedure detailed in Algorithm 1. As shown, DNO, similar to LGD Song et al. (2023a), is tuning-free as it does not require tuning the network parameter $\theta$. However, it requires more time for generation compared to direct sampling, as the optimization is integrated with the sampling process, meaning the optimization is performed for each new sample generated. Despite this, as we will demonstrate in Section 5, the extra time needed for the DNO approach is an acceptable trade-off for obtaining high-reward samples.
Figure 3: Overview of the DNO procedure with the DDIM sampling algorithm: DNO seeks to optimize only those Gaussian noise vectors $\{x_{T},z_{1},z_{2}...,z_{T}\}$ to maximize the reward value of a single generated sample $x_{0}$. To facilitate the gradient backpropagation from $x_{0}$ to $\{x_{T},z_{1},z_{2}...,z_{T}\}$, we leverage the technique of gradient checkpointing. It is worth noting that when using $\eta=0$ for DDIM sampling, there is no need to compute the gradient for $z_{1},...,z_{T}$, as the generated sample $x_{0}$ depends exclusively on $x_{T}$. When computing the gradient from $r(x_{0})$ to $x_{0}$, we can use either ground-truth gradient $\nabla r$ or an estimated gradient $\widehat{\nabla r}$, depending on whether the reward function $r(\cdot)$ is differentiable.
@@
While the concept of DNO has been proposed in previous works, many of its technical details remain insufficiently explored. On one hand, there is a lack of a comprehensive framework concerning the design choices, theoretical understanding, and practical challenges associated with DNO for aligning diffusion models. On the other hand, the field has yet to systematically investigate whether DNO, as a tuning-free method for aligning diffusion models, can achieve competitive performance compared to tuning-based methods. In this work, we aim to conduct a thorough study on DNO.
In the following two sections, we dive deep to understand the theoretical foundation of DNO and discuss a critical design choice ignored in previous works.
2.1 Understanding Noise Optimization
Example 1. First, we present a simple example to visualize the process of noise optimization. Specifically, we trained a toy diffusion model for generating uniform distribution on a ring with a radius between $0.8$ and $1.2$, and the initial distribution is visualized in Figure 4(a), where each red point denotes a single sample drawn from the trained diffusion model. We then solve the noise optimization problem (8) with the reward function $r(x)=\sin(4\pi x[1])+\sin(4\pi x[2])-\left((x[1]-1)^{2}-x[2]^{2}\right)/5$, a highly nonconvex function with many local maxima. To perform this optimization, we solve the DNO problem (8) using gradient ascent with the learning rate set to 0.01. In Figures 4(b), 4(c), and 4(d), we visualize the optimized samples after 10, 50, and 100 gradient steps, respectively. We can observe that the distribution of the samples shifts toward a distribution on the local maxima of the reward function.
@@ 2.1 Understanding Noise Optimization
Figure 4: Example 1: Evolution of the sample distribution of a toy diffusion model while running DNO to maximize a non-convex reward function.
Noise Optimization as Sampling from an Improved Distribution. As is evident from Figure 4, optimizing the noise vectors for diffusion models indeed leads to an improved distribution in terms of the reward scores. To rigorously describe this evolving process, we define an operator function $g$ to represent a single gradient step, i.e., $g(z)\stackrel{{\scriptstyle\text{def}}}{{=}}z+\ell\cdot\nabla_{z}r\circ M_{\theta}(z)$, where $\circ$ denotes the function composition operator and $\ell$ denotes the step size for gradient ascent. Additionally, we define the operator $g_{t}$, which denotes applying the gradient ascent step for $t$ steps, i.e., $g_{t}(z)=g(g_{t-1}(z))$ with $g_{0}$ being the identity mapping. With these notations, we can now express the distribution after $t$ gradient steps as $p_{t}(x)$, which is the distribution of $M_{\theta}(g_{t}(z))$ with $z\sim\mathcal{N}(0,I)$. In the following theorem, we demonstrate that there is a rigorous improvement after every single gradient step, i.e., the distribution $p_{t+1}(x)$ is provably better than $p_{t}(x)$ in terms of expected reward.
@@
As described in (11), provided that the step size $\ell$ is less than $\frac{2}{L}$, the distribution $p_{t+1}(x)$ is strictly better than the previous distribution $p_{t}(x)$ in terms of expected reward, as long as the second term is non-zero. Based on this result, it is natural to ask, The next question is: When does the distribution stop improving? Namely, when does the second term in (11) become zero. To answer this question, we provide a detailed discussion on this point and the proof for Theorem 1 in Appendix B. As described in (11), provided that the step size $\ell$ is less than $\frac{2}{L}$, the distribution $p_{t+1}(x)$ is strictly better than the previous distribution $p_{t}(x)$ in terms of expected reward, as long as the second term is non-zero. Based on this result, it is natural to ask: When does the distribution stop improving? Namely, when does the second term in (11) become zero. To answer this question, in Appendix B we provide a detailed discussion, together with the proof for Theorem 1.
@@
As previously introduced, there are two primary methods for sampling from pretrained diffusion models: one based on solving the ODE (2) and the other on solving the SDE (3). A critical difference lies in the fact that ODE sampling depends exclusively on the initial noise $x_{T}$, whereas SDE sampling is additionally influenced by the noise ${z_{t}}$ added at every step of the generation process. It has been noted that existing works on noise optimization (Ben-Hamu et al., 2024; Novack et al., 2024; Karunratanakul et al., 2023) have mainly concentrated on optimizing the initial noise $x_{T}$ for the ODE sampler.
Figure 5: ODE vs. SDE for optimization. The x-axis refers to the number of gradient ascent steps during optimization.
Different from preceding studies, we explore the utilization of the SDE sampler for noise optimization. Specifically, we employ the DDIM with $\eta=1$ (Song et al., 2020a) as the SDE sampler and propose to optimize both the added noise $z_{t}$ at every timestep and the initial noise $x_{T}$. In this context, the dimensionality of the optimized noise significantly surpasses that of ODE sampler, typically $T\cdot d$ v.s. $d$, where $d$ is the dimension for the learned distribution and $T$ is the number of discretization steps in the sampling process. With this higher dimension for optimization, we have empirically observed that optimizing the SDE-based generation process can be significantly faster than its ODE-based counterpart. This is illustrated in Figure 5, where we juxtapose the optimization speeds between the ODE (DDIM with $\eta=0$) and SDE samplers using both the Example 1 depicted in Figure 4 (Left) and optimizing for stable diffusion in alignment with the aesthetic reward (Right), a major experiment detailed in the subsequent section 5.2.
Understanding the Faster Speed of DNO with SDE-Based Sampling. Intuitively, the acceleration in optimization speed can be attributed to the finer-grained control over the generation process afforded by the SDE sampler compared to the ODE sampler. To formally state this intuition, we revisit the DDIM sampling algorithm in Algorithm 1. We consider the procedure of SDE-based sampling algorithm, DDIM with $\eta=1$, as defining the noise-to-sample mapping $M_{\theta}(x_{T},z_{1},\ldots,z_{T})$. An important observation is that the ODE sampling algorithm, DDIM with $\eta=0$, can also be expressed using the same $M_{\theta}(x_{T},z_{1},\ldots,z_{T})$, with the distinction that the noise $z_{1},\ldots,z_{T}$ becomes deterministic and dependent on $x_{T}$, rather than sampled from a Gaussian distribution. Specifically, if we define the deterministic noise vectors as:
$z_{t}^{\text{ODE}}\stackrel{{\scriptstyle\text{def.}}}{{=}}\frac{\sqrt{1-\alpha_{t-1}}-\sqrt{1-\alpha_{t-1}-\sigma_{t}^{2}}}{\sigma_{t}}\epsilon_{\theta}(x_{t},t),\quad\text{for }t=1,\ldots,T,$ (12)
where $\sigma_{t}=\sqrt{(1-\alpha_{t-1}/(1-\alpha_{t})}\sqrt{1-\alpha_{t}/\alpha_{t-1}}$, then the sampling process of DDIM with $\eta=0$ can be expressed as $M_{\theta}(x_{T},z_{1}^{\text{ODE}},\ldots,z_{T}^{\text{ODE}})$. In this context, the advantage of SDE-based sampling becomes evident:
$\underbrace{\max_{x_{T},z_{1},\ldots,z_{T}}r\left(M_{\theta}(x_{T},z_{1},\ldots,z_{T})\right)}_{\text{DNO with SDE}}\geq\underbrace{\max_{x_{T}}r\left(M_{\theta}(x_{T},z_{1}^{\text{ODE}},\ldots,z_{T}^{\text{ODE}})\right)}_{\text{DNO with ODE}},$ (13)
meaning that running DNO with SDE will yield better results, or at least as good as DNO with ODE for aligning diffusion models.
Based on this conclusion, our work will focus on optimizing the SDE-based sampling (DDIM with $\eta=1$) for the remainder of the study. Additionally, we fix the number of generation steps $T$ to 50 throughout this work for simplicity.